SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (April 25, 2006)

Attendees: Rainer Bartoldus, Al Eisner, Fang Fang, Selina Li, Dong Su

Operations

  • There was an unscheduled downtime to work on the LER overheated kicker on April 21st . We used this opportunity to test the TSF firmware (see the DCT/DCZ section below).
  • PEP might have a scheduled downtime for a chamber replacement next Monday which will last probably two days.

DCT/DCZ

  • Marc was able to reproduce the user reset problem on his test board, and with recommendation from Xilinx, he compiled a new version of the TSF firmware. Selina tested this latest version at the teststand and IR2. All went well except that there were ZPD damages (TSF -> ZPD frame bit errors) so we rolled back to an older version of the firmware. See HN posted and its follow-up postings.
  • After some discussions of the ZPD damages, we decided to implement a L3 filter for ZPD damages to prevent damaged events getting into our physics data stream. Level 3 already treats ZPD damage properly, by passing any damaged event it can find (up to a rate limit). The offline Event Filter has to do the opposite thing and keep the events from passing. It currently detects but ignores the BLT/ZPD damage. What's involved in keeping the ZPD-damaged events out of the physics collections is a new Filter configuration that tells it to redirect these along with other damage. Rainer will decide with the Computing Management from which point in time we want to activate that new configuration. For all future run, since just before the damaged run, or retroactively since Run 5.
  • Fang has been looking at the events with out-of-order errors seen in the ROM logfile. It looks like the events with out-of-order also have out-of-synch errors as well. She showed us a logfile that contains events with the errors indicating information such as crate, board number, header, etc. that might tell us more info about these events. Su Dong suggested that we should unpack the word to get the DAQ buffer number, tag, and count.
  • The out-of-order error showed up again in the TSF-X ROM logfile on April 23rd. It looks like it might be correlated with the fact that we are using the version of the TSF firmware which showed TSF-X having out-of-order errors every once a while in the past.

Online

  • Fang will put in a delay of 2 (that is 2/60 seconds) in the DCZ ROM code after the user reset even thought the current version of the firmware doesn't really require it.
  • Fang tested the DCZ lib after the removal of the teststand dependency during the MD at IR2 today.
  • Soon we will have a new online build based on 18.6.5 which contains changes in the L3 configuration that works with ROOT, fixed bug for high occupancy events, extra monitoring plots, etc..

GLT

  • We took a run 63994 with the configuration key "PHYSICS_TIGHT" last Tuesday. See HN posted. Rained showed the resulting reduction in the L1 Trigger rate in the detector operation HN posted. We decided that it's safe and beneficial to use this new config to take data, so later today Rainer will update the trig config to use as the default physics key.
  • Su Dong is making some projections for the trigger rate/configuration for David MacFarlane to present at the DOE B Factory operations review. The big question is if there is a way trigger can handle the projected high luminosity; that means we need to be albe to reduce the L1 trigger rate from 7kHz to the experimental limit of ~5kHz. The MC physics samples Eric generated last year don't contain MC true information, so it's a bit difficult to study the trigger configuration.

Data Quality and Monitoring

  • Al pointed out that the error bar issue still needs to be addressed for the OPR lumi ratio strip chart.
  • Al found out from the DQG meeting that there is a way to group the data by integrated luminosity for use in luminosity ratio strip chart.
  • For next week's TRG operation meeting, we will be discussing trickle injection monitoring.
  • Please perform the OPR data quality check for runs 64100-64104 (inclusive) very carefully since there are the runs taken with the latest version of the TSF firmware and there are ~5% ZPD damages in them.
  • There is evidence showing that the hadron ratio depends on background (presented at the BaBar/PEP meeting). It seems that the ratio goes up when background is worse. This might be due to background from proton.
  • Rainer also pointed out that the offline lumi is not right. See HN posted. The Friday AWG will discuss this more in detail.
  • Rahmat updated the JAS page for DCT->GLT and EMT->GLT transmission error alarm threshold for emailing and paging experts. This is to reflect the expected growth of error rate with the increase of luminosity.
  • It was suggested that we add the GLT->FCT consistency check as part of the L1T fast monitoring as well.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.